OpenStudy (anonymous):

what is the approxomite circumference of a circle with a radius of 30 inches. Use 3.14 = pi reward will Be GIVEN, HELP ME PLEASE A.47.1 B.94.2 C.62.8 D.188.4

jimthompson5910 (jim_thompson5910):

Use the formula C = 2*pi*r where r = radius

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what do you get

OpenStudy (anonymous):

94.2

jimthompson5910 (jim_thompson5910):

C = 2*pi*r C = 2*3.14*r ... replace pi with 3.14 C = 2*3.14*30 ... replace r with 30 C = ???

OpenStudy (anonymous):

188.4

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

1 more

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

ok, hold on

OpenStudy (anonymous):

a rectangle has an area of 25 square centimeters and a length of 10 centimeters. What is the width of the rectangle? A.2.5 B.7.5 C.15 D.35

jimthompson5910 (jim_thompson5910):

A = LW 25 = 10*W solve for W

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

divide both sides by 10 to isolate W

OpenStudy (anonymous):

2.5???? idk

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (anonymous):

hurray, one more?

OpenStudy (anonymous):

wat is the appoxomite area of the fiqure?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

is this a circle? or an oval? if possible, please post a screenshot of the problem

OpenStudy (anonymous):

oval

jimthompson5910 (jim_thompson5910):

can you post a screenshot?

OpenStudy (anonymous):

no srry

jimthompson5910 (jim_thompson5910):

I'm not sure exactly if this is the exact figure (and if the dimensions are correct)